_____24. A prismatoid has an upper base of 90% , a lower base of 2100% , an altitude of 100 cm, and a midsection of 120? . What is its volume? (with solution)

a) 7000 cm3 b) 11 500 cm3 c) 13 000 cm3 d) 17 500 cm3
1
Expert's answer
2020-12-13T19:24:36-0500
"V=\dfrac{L}{6}(A_1+4A_m+A_2)"

where

"A_1" and "A_2=" areas of parallel bases

"A_m=" area of the section midway between "A_1" and "A_2"

"L=" the altitude



"V=\dfrac{100}{6}(90+4(120)+210)=13000(cm^3)"

c) 13 000 cm3